how do you work out such an equation? x-2+3(x-4)=2(x-2)
Answer by Theo(13342)

x-2+3(x-4)=2(x-2)

simplify the equation by removing parentheses.

you get:

x - 2 + 3*x - 12 = 2*x - 4

combine like terms to get:

4*x - 14 = 2*x - 4

subtract 2*x from both sides of the equation to get:

2*x - 14 = -4

add 14 to both sides of the equation to get:

2*x = 10

divide both sides of the equation by 2 to get:

x = 5

substitute 5 for x in your original equation to see if it holds true.

your original equation is:

x-2+3(x-4)=2(x-2)

this becomes:

5 - 2 + 3(5-4) = 2(5-2) which becomes:

5 - 2 + 3(1) = 2(3) which becomes:

5 - 2 + 3 = 6 which becomes:

6 = 6 which is true so the value of 5 for x is good.